Write each conic sectioning standard form. Identify the conic section. Graph and label the appropriate parts.
1. 5x^2-y=0
2. x^2+4x+y^2+4y-8=0

Answer by richwmiller(17219)

5x^2-y=0 is a parabola
y=5x^2
x^2+4x+y^2+4y-8=0 is a circle
(x+2)^2+(y+2)^2=16